# KiDS-Legacy: Consistency of cosmic shear measurements and joint cosmological constraints with external probes

### v1

@@
11institutetext: Ruhr University Bochum, Faculty of Physics and Astronomy, Astronomical Institute (AIRUB), German Centre for Cosmological Lensing, 44780 Bochum, Germany 22institutetext: School of Mathematics, Statistics and Physics, Newcastle University, Herschel Building, NE1 7RU, Newcastle-upon-Tyne, United Kingdom 33institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, United Kingdom 44institutetext: Leiden Observatory, Leiden University, P.O.Box 9513, 2300RA Leiden, The Netherlands 55institutetext: Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, United Kingdom 66institutetext: Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 77institutetext: Waterloo Centre for Astrophysics, University of Waterloo, Waterloo, ON N2L 3G1, Canada 88institutetext: Department of Physics and Astronomy, University of Waterloo, Waterloo, ON N2L 3G1, Canada 99institutetext: Institute for Theoretical Physics, Utrecht University, Princetonplein 5, 3584CC Utrecht, The Netherlands 1010institutetext: Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Av. Complutense 40, E-28040 Madrid, Spain $1$institutetext: Institute of Cosmology & Gravitation, Dennis Sciama Building, University of Portsmouth, Portsmouth, PO1 3FX, United Kingdom 1212institutetext: Universität Innsbruck, Institut für Astro- und Teilchenphysik, Technikerstr. 25/8, 6020 Innsbruck, Austria 1313institutetext: Dipartimento di Fisica e Astronomia ”Augusto Righi” - Alma Mater Studiorum Università di Bologna, via Piero Gobetti 93/2, I-40129 Bologna, Italy 1414institutetext: Istituto Nazionale di Astrofisica (INAF) - Osservatorio di Astrofisica e Scienza dello Spazio (OAS), via Piero Gobetti 93/3, I-40129 Bologna, Italy 1515institutetext: Istituto Nazionale di Fisica Nucleare (INFN) - Sezione di Bologna, viale Berti Pichat 6/2, I-40127 Bologna, Italy 1616institutetext: INAF - Osservatorio Astronomico di Padova, via dell’Osservatorio 5, 35122 Padova, Italy 1717institutetext: Institute for Computational Cosmology, Ogden Centre for Fundament Physics - West, Department of Physics, Durham University, South Road, Durham DH1 3LE, United Kingdom 1818institutetext: Centre for Extragalactic Astronomy, Ogden Centre for Fundament Physics - West, Department of Physics, Durham University, South Road, Durham DH1 3LE, United Kingdom 1919institutetext: Center for Theoretical Physics, Polish Academy of Sciences, al. Lotników 32/46, 02-668 Warsaw, Poland 2020institutetext: Institute for Particle Physics and Astrophysics, ETH Zürich, Wolfgang-Pauli-Strasse 27, 8093 Zürich, Switzerland 2121institutetext: Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona), Spain $2$institutetext: Department of Physics, University of Oxford, Denys Wilkinson Building, Keble Road, Oxford OX1 3RH, United Kingdom 2323institutetext: Donostia International Physics Center, Manuel Lardizabal Ibilbidea, 4, 20018 Donostia, Gipuzkoa, Spain 2424institutetext: The Oskar Klein Centre, Department of Physics, Stockholm University, AlbaNova University Centre, SE-106 91 Stockholm, Sweden 2525institutetext: Imperial Centre for Inference and Cosmology (ICIC), Blackett Laboratory, Imperial College London, Prince Consort Road, London SW7 2AZ, United Kingdom 2626institutetext: Zentrum für Astronomie, Universität Heidelberg, Philosophenweg 12, D-69120 Heidelberg, Germany 2727institutetext: Institute for Theoretical Physics, Philosophenweg 16, D-69120 Heidelberg, Germany 2828institutetext: Department of Physics “E. Pancini” University of Naples Federico II C.U. di Monte Sant’Angelo Via Cintia, 21 ed. 6, 80126 Naples, Italy 2929institutetext: INAF – Osservatorio Astronomico di Capodimonte, Salita Moiariello 16, I-80131 Napoli, Italy 3030institutetext: INFN, Sez. di Napoli, via Cintia, 80126, Napoli, Italy
@@
Benjamin Stölzner stoelzner@astro.rub.de11    Angus H. Wright 11    Marika Asgari 22    Catherine Heymans 1133    Hendrik Hildebrandt 11    Henk Hoekstra 44    Benjamin Joachimi 55    Konrad Kuijken 44    Shun-Sheng Li 1144    Constance Mahony 2222232311    Robert Reischke 1166    Mijin Yoon 44    Maciej Bilicki 1919    Pierre Burger 778866    Nora Elisa Chisari 9944    Andrej Dvornik 11    Christos Georgiou 2121    Benjamin Giblin 33    Joachim Harnois-Déraps 22    Priyanka Jalan 1919    Anjitha John William 1919    Shahab Joudaki 10101111    Giorgio Francesco Lesci 13131414    Laila Linke 1212    Arthur Loureiro 24242525    Matteo Maturi 26262727    Lauro Moscardini 131314141515    Nicola R. Napolitano 282829293030    Lucas Porth 66    Mario Radovich 1616    Tilman Tröster 2020    Maximilian von Wietersheim-Kramsta 17171818    Anna Wittje 11    Ziang Yan 11    Yun-Hao Zhang 3344
(Received ; accepted )
We present a cosmic shear consistency analysis of the final data release from the Kilo-Degree Survey (KiDS-Legacy). By adopting three tiers of consistency metrics, we compare cosmological constraints between subsets of the KiDS-Legacy dataset split by redshift, angular scale, galaxy colour and spatial region. We also review a range of two-point cosmic shear statistics. With the data passing all our consistency metric tests, we demonstrate that KiDS-Legacy is the most internally consistent KiDS catalogue to date. In a joint cosmological analysis of KiDS-Legacy and DES Y3 cosmic shear, combined with data from the Pantheon+ Type Ia supernovae compilation and baryon acoustic oscillations from DESI Y1, we find constraints consistent with Planck measurements of the cosmic microwave background with $S_{8}\equiv\sigma_{8}\sqrt{\Omega_{\rm m}/0.3}=0.814^{+0.011}_{-0.012}$ and $\sigma_{8}=0.802^{+0.022}_{-0.018}$.
@@ 1 Introduction
In the current era of precision cosmology, weak lensing surveys have probed the standard $\Lambda$-cold dark matter ($\Lambda$CDM) cosmological model to unprecedented precision. The weak gravitational lensing effect causes distortions in the shape of galaxy images, known as cosmic shear. This allows mapping the distribution of gravitating matter along the line of sight, which is sensitive to the shape and amplitude of the matter power spectrum. Since its first detection (Kaiser et al., 2000; Wittman et al., 2000; Van Waerbeke et al., 2000; Bacon et al., 2000), cosmic shear has become a primary cosmological probe for imaging galaxy surveys. Recent analyses from the three current stage-III weak lensing surveys, the ESO Kilo-Degree Survey (KiDS; Kuijken et al., 2015), the Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016), and the Subaru Hyper Suprime Cam Subaru Strategic Program (HSC; Aihara et al., 2018), have showcased the potential of cosmic shear measurements as a probe of the cosmological model, making it a main science driver for upcoming galaxy surveys conducted with the Vera C. Rubin Observatory (Ivezić et al., 2019), the Euclid satellite (Euclid Collaboration et al., 2024), the Nancy Grace Roman Space Telescope (Spergel et al., 2015), and the Chinese Space Station Telescope (Gong et al., 2019).
Weak lensing studies mainly constrain the structure growth parameter $S_{8}=\sigma_{8}\sqrt{\Omega_{\rm m}/0.3}$, which combines the matter density parameter $\Omega_{\rm m}$ and the standard deviation of matter density perturbations in spheres of $8\,h^{-1}{\rm Mpc}$ radius, $\sigma_{8}$. Current cosmic shear measurements have yielded $S_{8}$ values that are lower (Asgari et al., 2021; Heymans et al., 2021; Abdalla et al., 2022; Amon et al., 2022; Secco et al., 2022; Dalal et al., 2023; Li et al., 2023b; Dark Energy Survey and Kilo-Degree Survey Collaboration et al., 2023) than values derived from observations of the cosmic microwave background (Planck Collaboration et al., 2020) at a $\sim 2\sigma$ level. However, there is no consensus about whether this feature, which is commonly referred to as the ?$S_{8}$ tension?, is a result of systematics in the data analysis or theoretical modelling, statistical fluctuations, or effects beyond the standard flat $\Lambda$CDM model.
@@
Given the unclear nature of the apparent $S_{8}$ tension, recent works have studied probes of the late Universe, finding consistency between independent cosmic shear surveys (Amon & Efstathiou, 2022; Amon et al., 2023; Longley et al., 2023; Dark Energy Survey and Kilo-Degree Survey Collaboration et al., 2023). In addition to tests of the consistency between independent data sets, tests of the internal consistency of a given data set are of particular importance in order to rule out systematic effects within one data set as the source of the inconsistency between different data sets (Efstathiou & Lemos, 2018; Köhlinger et al., 2019; Raveri et al., 2020; Li et al., 2021).
@@
This work is part of a series of KiDS-Legacy papers. The production process of all data products in the fifth KiDS data release (DR5), including shape measurements and the KiDS-Legacy sample selection, is presented in Wright et al. (2024). The calibration of the photometric redshift distribution is described in Wright et al. (2025a) and multi-band image simulations enabling a joint shear and redshift calibration are presented in Li et al. (2023a). The modelling of the covariance for the three main summary statistics is summarised in Reischke et al. (2024) and the angular clustering of KiDS-Legacy galaxies is analysed in Yan et al. (2025). The fiducial cosmic shear analysis is conducted in Wright et al. (2025b, hereafter W25). In this work, we perform several internal consistency tests of the KiDS-Legacy data, focusing on their impact on cosmological constraints inferred in the cosmic shear analysis. In particular, we split the KiDS-Legacy data set into various subsets based on redshift, spatial region, angular scale, and colour. We then perform a split cosmological analysis by modelling the observed data in each subset with a separate set of cosmological parameters. By evaluating several consistency metrics, we quantify the level of agreement between the data subsets. As established in the previous KiDS-1000 analysis (Asgari et al., 2021), constraints on cosmology are inferred from three different two-point statistics (COSEBIs, band powers, and two-point correlation functions). Here, we quantify the internal consistency between summary statistics.
@@
While current cosmic shear surveys cannot constrain both $\Omega_{\rm m}$ and $\sigma_{8}$ separately, the addition of external data allows for a breaking of the degeneracy. For this purpose we employ data from recent measurements of baryon acoustic oscillations (BAOs), redshift space distortions (RSDs), and Type Ia supernovae (SN Ia), which place tight constraints on the matter density. We perform a consistency test and a joint cosmological analysis of KiDS-Legacy data combined with data from the recent DESI Y1 BAO analysis (Adame et al., 2025), the earlier eBOSS (Alam et al., 2021) BAO and RSD analysis, and the Pantheon+ SN Ia compilation (Scolnic et al., 2022; Brout et al., 2022). Additionally, we conduct a joint analysis with DES Y3 cosmic shear data (Amon et al., 2022; Secco et al., 2022) and quantify the consistency between KiDS-Legacy and Planck CMB constraints (Planck Collaboration et al., 2020).
@@
This paper is structured as follows. In Sect. 2 we provide a brief summary of the KiDS-Legacy data and external data employed in this work. In Sect. 3 we review the theoretical model for weak lensing observables and discuss the metrics quantifying the internal consistency of the data. In Sect. 4 we present the results of the internal consistency tests. We provide the results of our joint cosmological constraints with external data in Sect. 5 and conclude in Sect. 6. Appendix A summarises the data properties of the KiDS-Legacy catalogue divided into sub-samples. In Appendix B we provide details about the estimate of the effective number of constrained parameters in our analysis.
@@ 2.1 KiDS-Legacy
The Kilo-Degree Survey (Kuijken et al., 2015; de Jong et al., 2015, 2017; Kuijken et al., 2019; Wright et al., 2024) is a public survey conducted by the European Southern Observatory with the VLT Survey Telescope (VST). The KiDS and the complementary VISTA Kilo-Degree Infrared Galaxy Survey (VIKING; Edge et al., 2013) combine optical and near-infrared imaging in nine photometric bands. In this work, we analyse weak lensing data from the fifth and final data release (DR5). Here, we provide a brief summary of the survey and refer the reader to Wright et al. (2024) for the details.
The KiDS-DR5 data set consists of imaging data covering an area of 1347 ${\rm deg}^{2}$ on sky divided into two distinct stripes across the celestial equator in the North Galactic Cap and across the South Galactic Pole, respectively. All sources are observed in four optical bands ($u,g,r$, and $i$) with the VST as well as in five near-infrared bands ($Z,Y,J,H$, and $K_{\rm s}$) from VIKING. In comparison to the previous data release, KiDS-DR5 features a second pass of $i$-band observations and an increase of $34\%$ in survey area. The lensing sample is obtained via a masking process, selecting sources with high-quality data in all photometric bands, and applying a sequence of cuts on magnitude, colour, and lensing-related quantities. This sample, dubbed KiDS-Legacy, contains approximately 43 million sources on $967\,{\rm deg}^{2}$ of sky, corresponding to an effective number density $n_{\rm eff}=8.79\,{\rm arcmin}^{-2}$.
Photometric redshift estimates of KiDS-Legacy sources are computed via the Bayesian Photo-z code (bpz; Benítez, 2000). The deeper $i$-band depth and a significantly larger spectroscopic calibration data set allow for a higher photometric redshift limit of $z_{\rm B}=2$ compared to previous KiDS analyses, enabling the addition of another redshift bin. The sources in the KiDS-Legacy lensing sample are divided into six approximately equi-populated bins via their photometric redshift estimates. The redshift distributions of the resulting bins are inferred via a direct calibration with deep spectroscopic surveys using self-organising maps (SOMs; Lima et al., 2008; Masters et al., 2015; Wright et al., 2020a). While this calibration method was utilised in earlier KiDS analyses (Wright et al., 2020b; Hildebrandt et al., 2021), the KiDS-Legacy redshift calibration features several improvements such as the use of one SOM per tomographic bin instead of one overall SOM, a selection of sources via SOM-derived gold weights, and accounting for prior volume effects. Furthermore, the SOM redshift distributions are calibrated with the multicolour SKiLLS simulations (Li et al., 2023a). For a detailed discussion of the improved calibration method, we refer the reader to the KiDS-Legacy redshift calibration manuscript (Wright et al., 2025a).

### v2

@@
11institutetext: Ruhr University Bochum, Faculty of Physics and Astronomy, Astronomical Institute (AIRUB), German Centre for Cosmological Lensing, 44780 Bochum, Germany 22institutetext: School of Mathematics, Statistics and Physics, Newcastle University, Herschel Building, NE1 7RU, Newcastle-upon-Tyne, United Kingdom 33institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, United Kingdom 44institutetext: Leiden Observatory, Leiden University, P.O.Box 9513, 2300RA Leiden, The Netherlands 55institutetext: Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, United Kingdom 66institutetext: Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 77institutetext: Waterloo Centre for Astrophysics, University of Waterloo, Waterloo, ON N2L 3G1, Canada 88institutetext: Department of Physics and Astronomy, University of Waterloo, Waterloo, ON N2L 3G1, Canada 99institutetext: Institute for Theoretical Physics, Utrecht University, Princetonplein 5, 3584CC Utrecht, The Netherlands 1010institutetext: Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Av. Complutense 40, E-28040 Madrid, Spain 1111institutetext: Institute of Cosmology & Gravitation, Dennis Sciama Building, University of Portsmouth, Portsmouth, PO1 3FX, United Kingdom 1212institutetext: Universität Innsbruck, Institut für Astro- und Teilchenphysik, Technikerstr. 25/8, 6020 Innsbruck, Austria 1313institutetext: Dipartimento di Fisica e Astronomia ”Augusto Righi” - Alma Mater Studiorum Università di Bologna, via Piero Gobetti 93/2, I-40129 Bologna, Italy 1414institutetext: Istituto Nazionale di Astrofisica (INAF) - Osservatorio di Astrofisica e Scienza dello Spazio (OAS), via Piero Gobetti 93/3, I-40129 Bologna, Italy 1515institutetext: Istituto Nazionale di Fisica Nucleare (INFN) - Sezione di Bologna, viale Berti Pichat 6/2, I-40127 Bologna, Italy 1616institutetext: INAF - Osservatorio Astronomico di Padova, via dell’Osservatorio 5, 35122 Padova, Italy 1717institutetext: Institute for Computational Cosmology, Ogden Centre for Fundament Physics - West, Department of Physics, Durham University, South Road, Durham DH1 3LE, United Kingdom 1818institutetext: Centre for Extragalactic Astronomy, Ogden Centre for Fundament Physics - West, Department of Physics, Durham University, South Road, Durham DH1 3LE, United Kingdom 1919institutetext: Center for Theoretical Physics, Polish Academy of Sciences, al. Lotników 32/46, 02-668 Warsaw, Poland 2020institutetext: Institute for Particle Physics and Astrophysics, ETH Zürich, Wolfgang-Pauli-Strasse 27, 8093 Zürich, Switzerland 2121institutetext: Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona), Spain 2222institutetext: Department of Physics, University of Oxford, Denys Wilkinson Building, Keble Road, Oxford OX1 3RH, United Kingdom 2323institutetext: Donostia International Physics Center, Manuel Lardizabal Ibilbidea, 4, 20018 Donostia, Gipuzkoa, Spain 2424institutetext: The Oskar Klein Centre, Department of Physics, Stockholm University, AlbaNova University Centre, SE-106 91 Stockholm, Sweden 2525institutetext: Imperial Centre for Inference and Cosmology (ICIC), Blackett Laboratory, Imperial College London, Prince Consort Road, London SW7 2AZ, United Kingdom 2626institutetext: Zentrum für Astronomie, Universität Heidelberg, Philosophenweg 12, D-69120 Heidelberg, Germany 2727institutetext: Institute for Theoretical Physics, Philosophenweg 16, D-69120 Heidelberg, Germany 2828institutetext: Department of Physics “E. Pancini” University of Naples Federico II C.U. di Monte Sant’Angelo Via Cintia, 21 ed. 6, 80126 Naples, Italy 2929institutetext: INAF – Osservatorio Astronomico di Capodimonte, Salita Moiariello 16, I-80131 Napoli, Italy 3030institutetext: INFN, Sez. di Napoli, via Cintia, 80126, Napoli, Italy 3131institutetext: Kapteyn Institute, University of Groningen, PO Box 800, NL 9700 AV Groningen
@@
Benjamin Stölzner stoelzner@astro.rub.de    Angus H. Wright    Marika Asgari    Catherine Heymans    Hendrik Hildebrandt    Henk Hoekstra    Benjamin Joachimi    Konrad Kuijken    Shun-Sheng Li    Constance Mahony    Robert Reischke    Mijin Yoon    Maciej Bilicki    Pierre Burger    Nora Elisa Chisari    Andrej Dvornik    Christos Georgiou    Benjamin Giblin    Joachim Harnois-Déraps    Priyanka Jalan    Anjitha John William    Shahab Joudaki    Giorgio Francesco Lesci    Laila Linke    Arthur Loureiro    Matteo Maturi    Lauro Moscardini    Nicola R. Napolitano    Lucas Porth    Mario Radovich    Tilman Tröster    Edwin Valentijn    Maximilian von Wietersheim-Kramsta    Anna Wittje    Ziang Yan    Yun-Hao Zhang
(Received 31 March 2025 / Accepted 31 July 2025)
We present a cosmic shear consistency analysis of the final data release from the Kilo-Degree Survey (KiDS-Legacy). By adopting three tiers of consistency metrics, we compared cosmological constraints between subsets of the KiDS-Legacy dataset split by redshift, angular scale, galaxy colour, and spatial region. We also reviewed a range of two-point cosmic shear statistics. As all the data passed our set of consistency metric tests, we demonstrate that KiDS-Legacy is the most internally consistent KiDS catalogue to date. In a joint cosmological analysis of KiDS-Legacy and DES Y3 cosmic shear, combined with data from the Pantheon+ Type Ia supernovae compilation and baryon acoustic oscillations from DESI Y1, we report constraints that are consistent with Planck measurements of the cosmic microwave background, with $S_{8}\equiv\sigma_{8}\sqrt{\Omega_{\rm m}/0.3}=0.814^{+0.011}_{-0.012}$ and $\sigma_{8}=0.802^{+0.022}_{-0.018}$.
@@ 1 Introduction
[…]
@@
Given the unclear nature of the apparent $S_{8}$ tension, recent works have studied probes of the late Universe, reporting a consistency between independent cosmic shear surveys (Amon & Efstathiou, 2022; Amon et al., 2023; Longley et al., 2023; Dark Energy Survey and Kilo-Degree Survey Collaboration et al., 2023). In addition to tests of the consistency between independent datasets, additional tests of the internal consistency of a given dataset are of particular importance to rule out systematic effects within one dataset as the source of the inconsistency between different datasets (Efstathiou & Lemos, 2018; Köhlinger et al., 2019; Raveri et al., 2020; Li et al., 2021).
@@
This work is part of a series of KiDS-Legacy papers. The production process of all data products in the fifth KiDS data release (DR5), including shape measurements and the KiDS-Legacy sample selection, is presented in Wright et al. (2024). The calibration of the photometric redshift distribution is described in Wright et al. (2025a) and multi-band image simulations enabling a joint shear and redshift calibration are presented in Li et al. (2023a). The modelling of the covariance for the three main summary statistics is summarised in Reischke et al. (2025) and the angular clustering of KiDS-Legacy galaxies is analysed in Yan et al. (2025). The fiducial cosmic shear analysis was conducted in Wright et al. (2025b, hereafter W25). In this work, we performed several internal consistency tests of the KiDS-Legacy data, focussing on their impact on cosmological constraints inferred in the cosmic shear analysis. In particular, we split the KiDS-Legacy dataset into various subsets based on redshift, spatial region, angular scale, and colour. We then performed a split cosmological analysis by modelling the observed data in each subset with a separate set of cosmological parameters. By evaluating several consistency metrics, we quantified the level of agreement between the data subsets. As established in the previous KiDS-1000 analysis (Asgari et al., 2021), constraints on cosmology were inferred from three different two-point statistics (COSEBIs, band powers, and two-point correlation functions). Here, we quantified the internal consistency between summary statistics.
@@
While current cosmic shear surveys cannot constrain both $\Omega_{\rm m}$ and $\sigma_{8}$ separately, the addition of external data allows for a breaking of the degeneracy. For this purpose, we employed data from recent measurements of baryon acoustic oscillations (BAOs), redshift space distortions (RSDs), and Type Ia supernovae (SN Ia), which place tight constraints on the matter density. We performed a consistency test and a joint cosmological analysis of KiDS-Legacy data combined with data from the recent DESI Y1 BAO analysis (Adame et al., 2025), the earlier eBOSS (Alam et al., 2021) BAO and RSD analysis, and the Pantheon+ SN Ia compilation (Scolnic et al., 2022; Brout et al., 2022). Additionally, we conducted a joint analysis with DES Y3 cosmic shear data (Amon et al., 2022; Secco et al., 2022) and quantified the consistency between KiDS-Legacy and Planck CMB constraints (Planck Collaboration et al., 2020).
@@
This paper is structured as follows. In Sect. 2, we provide a brief summary of the KiDS-Legacy data and external data employed in this work. In Sect. 3, we review the theoretical model for weak lensing observables and discuss the metrics quantifying the internal consistency of the data. In Sect. 4, we present the results of the internal consistency tests. We provide the results of our joint cosmological constraints with external data in Sect. 5 and present our conclusions in Sect. 6. Appendix A summarises the data properties of the KiDS-Legacy catalogue divided into sub-samples. In Appendix B, we provide details on our estimations of the effective number of constrained parameters in our analysis. Appendix C presents a sensitivity analysis of our consistency metrics.
@@ 2.1 KiDS-Legacy
The Kilo-Degree Survey (Kuijken et al., 2015; de Jong et al., 2015, 2017; Kuijken et al., 2019; Wright et al., 2024) is a public survey conducted by the European Southern Observatory with the VLT Survey Telescope (VST). KiDS and the complementary VISTA Kilo-Degree Infrared Galaxy Survey (VIKING; Edge et al., 2013) combine optical and near-infrared imaging in nine photometric bands. In this work, we analyse weak lensing data from the fifth and final data release (DR5). Here, we provide a brief summary of the survey and refer to Wright et al. (2024) for details.
The KiDS-DR5 dataset consists of imaging data covering an area of 1347 ${\rm deg}^{2}$ on-sky, divided into two distinct stripes across the celestial equator in the North Galactic Cap and across the South Galactic Pole, respectively. All sources were observed in four optical bands ($u,g,r$, and $i$) with the VST as well as in five near-infrared bands ($Z,Y,J,H$, and $K_{\rm s}$) from VIKING. In comparison to the previous data release, KiDS-DR5 features a second pass of $i$-band observations and an increase of $34\%$ in survey area. The lensing sample was obtained via a masking process, selecting sources with high-quality data in all photometric bands and applying a sequence of cuts on magnitude, colour, and lensing-related quantities. This sample, dubbed KiDS-Legacy, contains approximately 43 million sources on $967\,{\rm deg}^{2}$ of sky, corresponding to an effective number density of $n_{\rm eff}=8.79\,{\rm arcmin}^{-2}$.
Photometric redshift estimates of KiDS-Legacy sources were computed via the Bayesian Photo-z code (bpz; Benítez, 2000). The deeper $i$-band depth and a significantly larger spectroscopic calibration dataset allow for a higher photometric redshift limit of $z_{\rm B}=2$ compared to previous KiDS analyses, enabling the addition of another redshift bin. The sources in the KiDS-Legacy lensing sample were divided into six approximately equi-populated bins via their photometric redshift estimates. The redshift distributions of the resulting bins were inferred via a direct calibration with deep spectroscopic surveys using self-organising maps (SOMs; Lima et al., 2008; Masters et al., 2015; Wright et al., 2020a). While this calibration method was utilised in earlier KiDS analyses (Wright et al., 2020b; Hildebrandt et al., 2021), the KiDS-Legacy redshift calibration features several improvements such as the use of one SOM per tomographic bin instead of one overall SOM and a selection of sources via SOM-derived gold weights, while also accounting for prior volume effects. Furthermore, the SOM redshift distributions were calibrated with the multi-colour SKiLLS simulations (Li et al., 2023a). For a detailed discussion of the improved calibration method, we refer to the KiDS-Legacy redshift calibration manuscript (Wright et al., 2025a).